# Unsupervised Keypoints from Pretrained Diffusion Models

### v1

@@ Abstract
Unsupervised learning of keypoints and landmarks has seen significant progress with the help of modern neural network architectures, but performance is yet to match the supervised counterpart, making their practicability questionable. We leverage the emergent knowledge within text-to-image diffusion models, towards more robust unsupervised keypoints. Our core idea is to find text embeddings that would cause the generative model to consistently attend to compact regions in images (i.e. keypoints). To do so, we simply optimize the text embedding such that the cross-attention maps within the denoising network are localized as Gaussians with small standard deviations. We validate our performance on multiple datasets: the CelebA, CUB-200-2011, Tai-Chi-HD, DeepFashion, and Human3.6m datasets. We achieve significantly improved accuracy, sometimes even outperforming supervised ones, particularly for data that is non-aligned and less curated. Our code is publicly available and can be found through our project page: https://ubc-vision.github.io/StableKeypoints/.
@@ 3.1 Attention maps in diffusion networks
Specifically, in the transformer part of the model, denote $\Phi_{l}^{c}(\cdot)$ and $\Psi_{l}^{c}(\cdot)$ as the $c$-th head and the $l$-th linear layers of the U-Net. We calculates the query as $\mathbf{Q}_{l}^{c}=\Phi_{l}^{c}(\mathbf{z}_{t{=}1})\in\mathbb{R}^{(H\times W)\times D_{l}}$, [^1] We choose $t{=}1$ where $T{=}50$ steps via hyper-parameter tuning. and the key from the language embedding $\mathbf{K}_{l}^{c}=\Psi_{l}^{c}(\mathbf{e})\in\mathbb{R}^{N\times D_{l}}$, where $N$ is the number of tokens, $C$ the number of heads in the transformer attention layer, $H$ and $W$ are image height and width at that specific layer in U-Net, and $D_{l}$ the dimensionality of the layer. Given query and key, the cross-attention map $\mathbf{M}_{l}\in\mathbb{R}^{(H\times W)\times N}$ is then computed by applying softmax along the $N$ dimension, and average pooling along the $C$ dimension:
@@ 3.2 Optimizing to find the keypoint embeddings
We first start by treating the attention map $\tilde{\mathbf{M}}_{n}$ as a probability map of where the associated embedding $\mathbf{e}_{n}$ is located by applying the softmax function. We write
$\mathbf{P}_{n}=\exp(\tilde{\mathbf{M}}_{n})\>/\>\sum_{h,w}(\exp(\tilde{\mathbf{M}}_{n}))\in\mathbb{R}^{H\times W},$ (6)
where we encourage localization by forcing $\mathbf{P}_{n}$ to be a single-mode Gaussian distribution located at its maximum. In more details, denoting the Gaussian image that shares the same maximum as $\mathbf{P}_{n}$ as $\mathbf{G}_{n}$, we write
$\mathcal{L}_{\text{localize}}=\mathbb{E}_{n}\|\mathbf{P}_{n}-\mathbf{G}_{n}\|^{2}.$ (7)
@@
$\mathbf{G}_{n}=\frac{\tilde{\mathbf{G}}_{n}}{\sum_{h,w}(\tilde{\mathbf{G}}_{n})},\>\>\tilde{\mathbf{G}}_{n}=\exp\left(-\frac{\|\mathbf{XY_{coord}}-\bm{\mu}_{n}\|_{2}^{2}}{2\sigma^{2}}\right),$ (9)
where $\mathbf{XY_{coord}}$ is a tensor of image coordinates, and normalization ensures the Gaussian target image can be interpreted as a probability distribution, compatible with $\mathbf{P}_{n}$.
@@
It is important to note that while $\mathcal{L}_{\text{localize}}$ in 7 at first glance seem to only encourage localization, it also enforces $\mathbf{P}_{n}$ to be mutually exclusive for different $n$ because of the softmax operation in 2. Should multiple embeddings become similar, their attention responses in 2 $\mathbf{Q}_{l}^{c}\cdot\mathbf{K}_{l}$ will become similar, resulting in the softmax of the attention map being a flat response (i.e. deviating from a Gaussian shape). In other words, 7 naturally enforces exclusivity with the help of 2.
@@
$\mathcal{N}(\kappa)=\text{ArgTop}_{\kappa}\>\>\{-\text{KL}(\mathbf{G}_{n},\mathbf{P}_{n})\}_{n=1}^{N}.$ (10)
@@ 4.1 Experimental setup
We note here that each dataset comes with its own metric. To make results more comparable across the human pose datasets, we report both their original metrics as well as the $\ell_{2}$ error when normalizing the image resolution to 128$\times$128.
@@
To evaluate the quality of unsupervised keypoints, one must relate them with human-annotated landmarks. Consistent with prior research [55], we use linear regression (without bias) to relate between unsupervised keypoints and human-annotated landmarks.
@@ 4.2 Experimental results
We report our results for each dataset in Tabs. 1, 2 and 3. As shown, except for the case when data is heavily processed and aligned (CelebA aligned in Tab. 1, CUB-aligned in Tab. 2, and Human 3.6M in Tab. 3), our method significantly outperforms the state of the art. The most visible gains are for the Tai-Chi-HD dataset, the most challenging among human pose datasets, and on CUB unaligned datasets. Note that for the CUB dataset and the Tai-Chi-HD datasets, we outperform even those that have been supervised with silhouettes or saliency maps.
@@
Finally, also note that for CUB-001, CUB-002, and CUB-003, these datasets are small. These datasets are non-aligned, have a large variability between individual images, and only contain 30 images each in the training set. Surprisingly, our method, just from 30 images, successfully identifies keypoints. These results highlight the potential of leveraging emergent (prior) knowledge within Stable Diffusion [44].
@@ 4.3 Ablation study
Figure 3: Generalization – we apply our learned text tokens (keypoints) to images that are from other datasets, including those that are of completely different domains. Our tokens generalize well for data of similar type, and surprisingly well even for some extreme cases.
Number of training images.
(a) CelebA dataset keypoints
(b) CUB-200-2011 dataset keypoints
(c) Tai-Chi-HD dataset keypoints
(d) DeepFashion dataset keypoints
(e) Human 3.6M dataset keypoints
(f) Unaligned Human 3.6M dataset keypoints
@@ 4.4 Generalization
We further test the generalization capacity of our learned keypoints. As they are effectively text embeddings, we can simply apply them to any image, including those completely outside of the training domain. We show qualitative examples in Fig. 3. As shown, even when applied to different datasets, they look reasonable. For example, in Fig. 2(a), when applying Tai-Chi-HD tokens to Human 3.6M, the tokens respond to the same locations on the human body as in Tai-Chi-HD. Surprising was when we applied CUB-200-2011 tokens to Tai-Chi-HD in Fig. 2(b)—they still responded to the body of the human being, reasonably consistently, although these tokens were trained to respond to birds. Of note are tokens two and six, which correspond to the front and back of the bird heads in Fig. 4—they also reply to the front and back of human heads. Applying CelebA tokens to Tai-Chi-HD in Fig. 2(c) also shows interesting outcomes, as tokens generally respond to human faces, despite the scale being drastically different between the two datasets. Finally, applying CelebA tokens to the CUB-200-2011 dataset in Fig. 2(d) shows mixed results—when it is ‘successful’ it focuses also on the faces of the bird, when it fails, it fails completely. These results hint that the keypoints (tokens) we have learned carry semantic meanings, as expected. We note that none of the baselines that we compare against are able to generalize beyond the dataset they were trained for. For a more comprehensive analysis please visit our project page: https://ubc-vision.github.io/StableKeypoints/.

### v3

@@ Abstract
Unsupervised learning of keypoints and landmarks has seen significant progress with the help of modern neural network architectures, but performance is yet to match the supervised counterpart, making their practicability questionable. We leverage the emergent knowledge within text-to-image diffusion models, towards more robust unsupervised keypoints. Our core idea is to find text embeddings that would cause the generative model to consistently attend to compact regions in images (i.e. keypoints). To do so, we simply optimize the text embedding such that the cross-attention maps within the denoising network are localized as Gaussians with small standard deviations. We validate our performance on multiple datasets: the CelebA, CUB-200-2011, Tai-Chi-HD, DeepFashion, and Human3.6m datasets. We achieve significantly improved accuracy, sometimes even outperforming supervised ones, particularly for data that is non-aligned and less curated. Our code is publicly available at the project page.
@@ 3.1 Attention maps in diffusion networks
Specifically, in the transformer part of the model, denote $\Phi_{l}^{c}(\cdot)$ and $\Psi_{l}^{c}(\cdot)$ as the $c$-th head and the $l$-th linear layers of the U-Net. We calculate the query as $\mathbf{Q}_{l}^{c}=\Phi_{l}^{c}(\mathbf{z}_{t{=}1})\in\mathbb{R}^{(H\times W)\times D_{l}}$, [^1] We choose $t{=}1$ where $T{=}50$ steps via hyper-parameter tuning. and the key from the language embedding $\mathbf{K}_{l}^{c}=\Psi_{l}^{c}(\mathbf{e})\in\mathbb{R}^{N\times D_{l}}$, where $N$ is the number of tokens, $C$ the number of heads in the transformer attention layer, $H$ and $W$ are image height and width at that specific layer in U-Net, and $D_{l}$ the dimensionality of the layer. Given query and key, the cross-attention map $\mathbf{M}_{l}\in\mathbb{R}^{(H\times W)\times N}$ is then computed via softmax along the $N$ dimension, and average pooling along the $C$ dimension:
@@ 3.2 Optimizing to find the keypoint embeddings
We encourage localization by forcing $\tilde{\mathbf{M}}_{n}$ to be a single-mode Gaussian distribution located at its maximum. In more details, denoting the Gaussian image that shares the same maximum as $\tilde{\mathbf{M}}_{n}$ as $\mathbf{G}_{n}$, we write
$\mathcal{L}_{\text{localize}}=\mathbb{E}_{n}\|\tilde{\mathbf{M}}_{n}-\mathbf{G}_{n}\|^{2}.$ (6)
@@
$\mathbf{G}_{n}=\exp\left(-\frac{\|\mathbf{XY_{coord}}-\bm{\mu}_{n}\|_{2}^{2}}{2\sigma^{2}}\right),$ (8)
where $\mathbf{XY_{coord}}$ is a tensor of image coordinates.
@@
It is important to note that while $\mathcal{L}_{\text{localize}}$ in 6 at first glance seem to only encourage localization, it also enforces $\tilde{\mathbf{M}}_{n}$ to be mutually exclusive for different $n$ because of the softmax operation in 2. Should multiple embeddings become similar, their attention responses in 2 $\mathbf{Q}_{l}^{c}\cdot\mathbf{K}_{l}$ will become similar, resulting in the softmax of the attention map being a flat response (i.e. deviating from a Gaussian shape). In other words, 6 naturally enforces exclusivity with the help of 2.
@@
$\mathcal{N}(\kappa)=\text{ArgTop}_{\kappa}\>\>\{-\text{KL}(\mathbf{G}_{n},\tilde{\mathbf{M}}_{n})\}_{n=1}^{N}.$ (9)
@@ 4.1 Experimental setup
Note that each dataset comes with its own metric. To make results more comparable across the human pose datasets, we report both their original metrics as well as the $\ell_{2}$ error when normalizing the image resolution to 128$\times$128.
@@
To evaluate the quality of unsupervised keypoints, one must relate them with human-annotated landmarks. As in prior research [55], we use linear regression (without bias) to relate between unsupervised keypoints and human-annotated landmarks.
@@ 4.2 Experimental results
(a) CelebA dataset keypoints
(b) CUB-200-2011 dataset keypoints
(c) Tai-Chi-HD dataset keypoints
(d) DeepFashion dataset keypoints
(e) Human 3.6M dataset keypoints
(f) Unaligned Human 3.6M dataset keypoints
@@
We report our results for each dataset in Tabs. 1, 2 and 3. As shown, except for the case when data is heavily processed and aligned (CelebA aligned in Tab. 1, CUB-aligned in Tab. 2, and Human 3.6M in Tab. 3), our method significantly outperforms the state of the art. The most visible gains are for the Tai-Chi-HD dataset, the most challenging among human pose datasets, and on CUB unaligned datasets. For the CUB dataset and the Tai-Chi-HD datasets, we outperform even those that have been supervised with silhouettes or saliency maps.
@@
Finally, also note that for CUB-001, CUB-002, and CUB-003, these datasets are small. These datasets are non-aligned, have a large variability between individual images, and only contain 30 images each in the training set. Our method, just from 30 images, successfully identifies keypoints. These results highlight the potential of leveraging emergent (prior) knowledge within Stable Diffusion [44].
@@ 4.3 Ablation study
Figure 4: Generalization – we apply our learned text tokens (keypoints) to images from other datasets, including those that are of completely different domains. Our tokens generalize well for data of similar type, and surprisingly well even for some extreme cases.
@@ 4.4 Generalization
We further test the generalization capacity of our learned keypoints. As they are effectively text embeddings, we can simply apply them to any image, including those completely outside of the training domain. We quantitatively evaluate our method and the previous best-performing method Autolink [13]. We find that even in these generalization experiments, our keypoints reach performance comparable to data-specific keypoints. Applying Tai-Chi-HD tokens to unaligned Human 3.6M achieves state-of-the-art performance despite using fewer keypoints (K=10 vs K=16). We also perform on par with the previous state of the art when we apply CUB-200-2011 tokens to Tai-Chi-HD—a case where the dataset gap is not only about the appearance but also beyond classes. Despite the drastic gap, our keypoints perform extremely well, leveraging the generalization power of large pre-trained diffusion models.
We show qualitative examples in Fig. 4. As shown, even when applied to different datasets, they look reasonable. For example, in Fig. 4(a), when applying Tai-Chi-HD tokens to Human 3.6M, the tokens respond to the same locations on the human body as in Tai-Chi-HD. Surprising was when we applied CUB-200-2011 tokens to Tai-Chi-HD in Fig. 4(b)—they still responded to the body of the human being, reasonably consistently, although these tokens were trained to respond to birds. Of note are tokens two and six, which correspond to the front and back of the bird heads in Fig. 3—they also reply to the front and back of human heads. Applying CelebA tokens to Tai-Chi-HD in Fig. 4(c) also shows interesting outcomes, as tokens generally respond to human faces, despite the scale being drastically different between the two datasets. Finally, applying CelebA tokens to the CUB-200-2011 dataset in Fig. 4(d) shows mixed results—when it is ‘successful’ it focuses also on the faces of the bird, when it fails, it fails completely. These results hint that the keypoints (tokens) we have learned carry semantic meanings, as expected. We note that none of the baselines that we compare against are able to generalize beyond the dataset they were trained for.
[…]
Ours 4.88 317.94 / 9.50 - / 8.6 18.60
Autolink [13] 16.92 535.61 / 16.00 - / 28.2 22.56
Table 5: Generalization – we quantitatively evaluate the performance of our keypoints on other datasets. Our Tai-Chi-HD keypoints applied to the unaligned Human3.6m setting reach state-of-the-art performance. Interestingly, our CUB keypoints applied to Tai-Chi-HD are on par with the previous state of the art, despite the differences between these datasets.